OpenStudy (anonymous):

Suppose that a droplet of mist is a perfect sphere and that, through condensation, the droplet picks up moisture at a rate proportional to its surface area. Show that under these circumstances the droplet's radius increases at a constant rate.

OpenStudy (jamesj):

Nice question. This means change in mass, delta m, is proportional to r^2.delta t. i.e., \[ dm = kr^2 \ dt \] for some constant k. Thus \[ \frac{dm}{dt} = kr^2 \] Now you want to show that \( dr/dt = constant \).

OpenStudy (anonymous):

does the d in "kr2 dt" mean delta? and where would I go from there? would i find the derivative of the surface area?

OpenStudy (jamesj):

the d is short hand for delta, yes. Now, as you know the volume of a sphere of radius r is \( (4/3)\pi r^3 \), hence is rho is the density of water, the mass of the sphere is \[ m = \frac{4\pi\rho}{3}r^3 \] and it follows that \[ \frac{dm}{dr} = (4\pi\rho)r^2 \] We now have an expression for \( dm/dr \) and for \( dm/dt \). Using the chain rule and/or whatever other techniques you like, find an expression for \( dr/dt \) and show it is constant.